1. Eight groups of cells are in parallel. Each group consists of  four cells, each of emf 4 V and internal resistance 0.5 Ώ, In series. Determine the power supplied to 4 Ώ resistor by this battery system
1
Expert's answer
2021-12-16T11:28:04-0500

Total emf of one group will be, "e = 4+4+4+4 = 16 V"

Total resistance of one group, "r = 0.5+0.5+0.5+0.5 = 2 \Omega"


Total emf of eight groups, "e = 16 V" as voltage remains same in parallel.

Total resistance of eight groups, "r' = \frac{2}{8} = 0.25 \Omega"


Total resistance of the circuit will be, "R = r' + 4 = 4+0.25 = 4.25\Omega"

Total current in the circuit is, "I = \frac{e}{R} = \frac{16}{4.25} = 3.77 A"


Power dissipated in 4 ohm resistance, "P = 4I^2 = 4 \times (3.77)^2 = 56.69 W"